1.Can you help me with this? ) Let’s assume that the combined population of Anchorage and Tampa was 25,000 in 1955. By 1960 the population of Anchorage had increased by 10% while Tampa had increased by 9%. If the total population increased by 2380 people, then what was the population of each city in 1955.
Thank you all

Answer by ankor@dixie-net.com(6567)

Let’s assume that the combined population of Anchorage and Tampa was 25,000
in 1955. By 1960 the population of Anchorage had increased by 10% while Tampa
had increased by 9%. If the total population increased by 2380 people,
then what was the population of each city in 1955.
:
Let x = population of Anchorage in 1955
then
(25000-x) = population of Tampa in 1955
:
Anchorage increase + Tampa increase = 2380 people
.10x + .09(25000-x) = 2380
.10x + 2250 - .09x = 2380
.10x - .09x = 2380 - 2250
.01x = 130
x =
x = 13,000 is the pop of Anchorage
and
25000 - 13000 = 12,000 is pop of Tampa
;
;
Check solution:
.10(13000) + .09(12000) =
1300 + 1080 = 2380